OpenStudy (anonymous):

When Θ = 5 pi over 6, what are the reference angle and the sign values for sine, cosine, and tangent? @ganeshie8

OpenStudy (mathmale):

Very glad to see you back on OpenStudy! Know in which quadrant you'll find this angle 5pi/6? Would you be willing to sketch this angle using the Draw feature?

OpenStudy (anonymous):

Nice to see you too! 5pi/6=2.617

OpenStudy (mathmale):

woohoo, would you find that angle in QI, QII, QIII or QIV?

OpenStudy (anonymous):

q1?

OpenStudy (mathmale):

Converting 5pi/6 to degrees would probably give you a clearer picture of this situation.

OpenStudy (mathmale):

\[\frac{ 5\pi }{ 6 }\frac{ 180 \deg }{ \pi }\]

OpenStudy (mathmale):

Please simplify that. 5pi/6 = how many degrees?

OpenStudy (anonymous):

150

OpenStudy (mathmale):

Right. Again, in which quadrant does this angle lie? This is important because it will tell you whether the adjacent side, the opposite side and the hypotenuse are positive or negative.

OpenStudy (anonymous):

4qudrant

OpenStudy (mathmale):



OpenStudy (mathmale):

Do you recognize this angle 5pi/6 as being equivalent to 150 deg? In what quadrant does 150 deg lie?

OpenStudy (anonymous):

in the 4th quadrant

OpenStudy (mathmale):

The quadrants are numbered 1, 2, 3, 4 beginning with the upper, right-hand one. Any angle between 0 and 90 deg is in Q1. Likewise, any angle between 90 and 180 deg is in Q2.

OpenStudy (anonymous):

oh so qdrt 2

OpenStudy (mathmale):

Please look at the diagram again. The vertical side of the triangle is positive. The horiz. side is negative. The hypotenuse, as always, is positive. OK? So, will the sine of 5pi/6 be pos. or neg., and why?

OpenStudy (anonymous):

postive idk

OpenStudy (anonymous):

hi mathmale its me can you please look into a question after you are finished with rhis one

OpenStudy (anonymous):

this

OpenStudy (mathmale):